OpenStudy (anonymous):

what is 6b+5-6b in siplest form

OpenStudy (anonymous):

Try grouping the terms with the 'b' variable and adding their coefficients.

OpenStudy (anonymous):

1-1=0 6b-6b=0 6b+5-6b=5

jhonyy9 (jhonyy9):

because 6b-6b=0 this will be 5

OpenStudy (anonymous):

thank guys